What is the pH of 0.0950 M Fe(NO3)3 (Ka of Fe3+ = 3.00x10-3)? Express your answer to two decimal places.
Answer #1

Fe³? = 0.0950 M

Now,

                 Fe³?    +    H?O    ------>    Fe(OH)²?    +    H?
IC:         0.0950                                       0                    0
C:              - x                                        + x                 + x
EC:     0.0950 – x                                    x                    x

Ka = x² / (0.0950-x)

3.00 x 10-3 = x² / (0.0950-x)

(3.00 x 10-3) (0.0950-x) = x²

x² + (3.00 x 10?³) x - (2.85 x 10??) = 0

By solving the quadratic equation, we have

x = 0.0154 and x = - 0.0184

Discarding the negative value of x, we have

x = 0.0154

So, [H?] = x = 0.0154 M

pH = - log [H?]

pH = - log (0.0154)

pH = 1.81